# MuSciClaims: Multimodal Scientific Claim Verification

### v1

@@ Abstract
Our results show most vision-language models are poor (0.3-0.5 F1), with even the best model only achieving 0.77 F1.
@@ 1 Introduction
However, scientific evidence is often presented as heterogeneous information-rich figures that support the important findings, claims and conclusions of experiments. Therefore, scientific claim verification requires both textual and visual understanding capabilities. To assess a claim, one needs to go over the figure and its caption, find the panel(s) with information relevant to the claim, combine this visual knowledge with textual information in the figure caption, and finally judging whether the claim is supported or not. While there is a large number of benchmarks on scientific figures, they focus on image captioning Hsu et al. (2021), question answering Kahou et al. (2017); Methani et al. (2020), or other reasoning tasks Yue et al. (2024a). There are no readily usable multimodal benchmarks for scientific claim verification. The closest work, ChartCheck Akhtar et al. (2024), poses a multimodal claim verification task but is restricted to simple data charts crawled from the web, which are substantially different from complex figures found in scientific articles.
To address this gap, we introduce
MuSciClaims 111We release the data at https://drive.google.com/file/d/1aSdnmxwAF2PJWoARByCjjcLKv6R2g7kF/view?usp=sharing., a multimodal benchmark for claim verification over figures in scientific literature.
We set forth two main desiderata for our benchmark: the dataset needs carefully constructed claims that are not supported or have contradictory information in the figures;
apart from quantifying model performance, the dataset should also be diagnostic in nature to identify specific model weaknesses.
Our dataset creation methodology is designed to meet these desiderata.
We first extract claims with inline references to figures from the results section of articles. We manually filter these to only retain claims that are clearly and unambiguously supported by the figures. Then, we create contradictory claims by perturbing these supporting claims. We devise a diverse set of perturbations that test specific capabilities required for claim verification including qualitative and quantitative reasoning, and epistemic observation-inference connections.
Last, we create a suite of diagnostic tasks associated with each claim to better understand model failures. Specifically, we design tasks that help uncover errors across aspects of basic visual understanding, evidence localization, cross-modal aggregation, and epistemic sensitivity. We ensure the integrity of the dataset through manual analysis. The resulting dataset consists of 918 (claim, figure) data points, equally balanced across 3 class labels, each accompanied by diagnostic questions.
@@
In fact, their basic visual understanding and cross-modal aggregation capabilities also need improvement.
@@ 2 Related Work
To create more complex QA benchmarks, ChartQA Masry et al. (2022) mixes 30k human and machine-generated questions; however the images are still limited to line, bar and pie charts. To cover more chart types, ArXivQA Li et al. (2024) extracts images with LLM-generated QA pairs from arXiv articles. MMC Liu et al. (2024) supports diverse tasks and chart types using free-form questions and open-ended answers.
Previous benchmarks rely heavily on chart annotations or table metadata as textual prompts to generate content, allowing models to easily obtain candidate answers while ignoring the charts’ visual logic. ChartBench Xu et al. (2023) includes both annotated and unannotated charts. While ChartX Xia et al. (2024) covers more chart types, its data and charts are synthesized and limited to ones that can be directly converted into a structural data format, e.g., CSV format. CharXiv Wang et al. (2024) consists of more than 2000 real-world charts with manually curated questions by human experts and answers validated by hand, panning 8 major subjects published on arXiv. The questions are either descriptive to understand basic chart data or reasoning-based to dig deeper into charts. MultiChartQA Zhu et al. (2025) is a benchmark designed to evaluate VLMs’ reasoning capabilities across multiple charts. However, the charts are not information-rich and no domain knowledge beyond what is stated in the charts is required to answer questions in these benchmarks.
In order to cover more image types, MMMU Yue et al. (2024a, b) collected multimodal questions from college exams, quizzes, and textbooks, covering six core disciplines, ranging from visual scenes like photographs and paintings to diagrams and tables, testing the perceptual capabilities of VLMs. While CURIE Cui et al. (2025) covers diverse scientific disciplines, its multimodal tasks are limited to biodiversity georeferencing and protein sequence reconstruction tasks. While answering them require domain knowledge, it doesn’t require a high level of expertise.
@@ 3 Creating
MuSciClaims contains a claim, its associated figure and caption and annotations about its relevant panels. Each claim, both original and perturbed, is also labeled with its relationship to the figure (Support, Neutral, Contradict). It also enables performing diagnostic tests. Models must identify the relevant panel as part of EvidenceLocalization and answer a question about the figure for BasicVisualUnderstanding. To perform Cross-ModalAggregation, model performance should drop when given either just the figure or the caption. For EpistemicSensitivity, model predictions must change across a pair of (original, perturbed) claims. Verifying whether a claim is supported by scientific evidence requires understanding different parts of the claim, locating and extracting information from multimodal sources, reasoning with it, and finally making a judgment.
@@
MuSciClaims, a dataset created from scientific articles in the life sciences.
We use peer-reviewed articles published in the Cell journal222We chose Cell since the articles are available for open access through https://www.cell.com. Also, the journal requires some uniformity in the source materials, which greatly simplifies the extraction process for our purposes., one of the highest impact factor venues in science.
@@ 3.1 Automatic Extraction
We extract figures and associated claims from the results section of the Cell journal articles, where the key findings and takeaways are described along with supporting evidence, often expressed within heterogeneous figures containing charts, microscopy images, or other diagrams.
@@ 3.2 Systematic Claim Perturbation
The claims extracted from the articles are grounded in the associated figures—i.e., the figures support333We rely on the scientific integrity of the published articles and assume that evidence support the asserted claims. the claims. To create an effective test bed, we also need claims that are not supported and have contradictory evidence in the figures. Further, we want to ensure that we test for a variety of reasons that could make a claim unsupported or contradictory with respect to the given evidence. To this end, we use a manual claim perturbation process to ensure meaningful perturbations, and ensure their quality through a second annotation process.
We manually analyzed the original claims to identify the main capabilities needed for checking a claim against the corresponding figures. Based on this analysis, we create four categories of perturbations: (i) Qualitative Inference—We replace directional terms with their opposites (e.g., “high concentration" to “low concentration"). This tests whether models can check if the asserted qualitative statements are supported via visual relationships between data points in a figure. (ii) Qualitative Relationship Inference —We edit comparisons (e.g., “X is stronger than Y" into “Y is stronger than X") to create the opposite conclusion from a figure. This checks for assessing qualitative relationships between variables via visual inference of similar relationships, (iii) Quantitative reasoning—We modify numerical values or ranges, primarily associated with experiment details such as statistical significance or experiment size to test for reasoning about the key quantities of interest. (iv) Epistemic Mismatch—This represents a disconnect between different forms of knowledge. We add perturbations that introduce inconsistencies between an observation that is visually true (i.e., supported by the figure) and its inference (which requires domain knowledge). This tests for the ability to carefully connect visually verified information with the textually asserted effect.
Our perturbations were created to ensure that the modified claim is a contradiction of the supported claim. This means that the figure which supports the original claim will, by extension, not support the modified claim. We verify the quality of the resulting perturbations through a second round of manual annotation. For a subset of data, three annotators were provided a supported claim as well as its perturbation. They are required to judge whether the perturbation contradicts the supported claim. We find that all three annotators agree that all perturbed claims are indeed contradictions of the corresponding supported claims (100% agreement).
@@ 3.4 Dataset Statistics
Through the process described above, we obtain 306 claims that are supported by a figure (Support), and 306 corresponding perturbed claims in contradiction to a given figure (Contradict).
@@
MuSciClaims contains 918 data points balanced equally across 3 class labels.
Along with this, each data point is annotated with panels most relevant to a claim, a question about the figure and information about perturbation types to support our diagnostic tests444We will release the data in accordance with the papers’ CC BY 4.0 license..
@@ 4.1 Evaluation Tasks
MuSciClaims also supports four other diagnostic tasks designed to assess a diverse set of capabilities required to effectively verify claims.
@@
We use Exact Match (EM) to judge whether a model answer is correct.
@@
(4) EpistemicSensitivity tests whether models consistently (and correctly) change their prediction across epistemic perturbation types of the same claim; they should predict support for the original and contradict for the perturbed claim.
@@
As part of §3.2, annotators also annotate whether they perturb the observation, the inference or both.
We devise a sensitivity metric to test for the same across (original, perturbed) claim pairs.
@@ 4.2 Models
We conduct the aforementioned evaluation on a set of different vision-language models (VLMs): gpt-4o-mini-2024-07-18 (4o-mini), gpt-4o-2024-11-20 (4o), claude-3-5-sonnet-20241022 (Sonnet), Phi-4 Multimodal Instruct (Phi-4), llava-v1.6-mistral-7b-hf (Llava-Next), Llama-3.2-11B Vision Instruct (Llama-3.2), Molmo-7B-D (Molmo), InternVL3-38B (InternVL3), Qwen2.5-VL-32B (Qwen2.5) and deepseek-VL2-small (DeepSeek).
@@
More details about models, prompts and setup are in Appendix A, F and D respectively.
@@ 5 Results
MuSciClaims.
@@
We observe that most models perform poorly on the task (D rows in Table 1), with overall F1 scores only ranging from 0.3-0.5. Only two models (out of ten) stand out: Sonnet (0.68 F1) and InternVL3 (0.74 F1). Per-class metrics also indicate that models perform well on Support and worst on Contradict.
Interestingly, we observe that models have high precision but low recall for Neutral and Contradict. This indicates that models cannot identify most of the claims in these categories but when they do identify them, they are correctly classified with low false positives. Contrastingly, models attain high recall and low precision on Support, indicating a bias towards assessing most of the claims as supported. Methods for better claim verification must alleviate such class biases.
@@
The RD rows in Table 1 represent results where models, given the figure and caption, first perform step-by-step reasoning on the claim and then generate their decision on the category of the claim. Results show that reasoning leads to improvements (0.02-0.04) for closed-source models and Llava-Next, but the gains are rather small. o4-mini, a model trained to analyze and do reasoning over images, achieves the highest performance.
@@ 6 Diagnostics Results
Going forward, we use our diagnostic tests to better understand the failure modes of 4o-mini, 4o, Sonnet and InternVL3.
@@
Finding the most relevant panels of figures is important to assess claims from information-rich figures. Table 2 shows how well models perform when prompted to first identify the associated panels, reason over them and make a decision (IRD), thereby testing EvidenceLocalization. Model performance deteriorates when localizing before reasoning (IRD) as compared to reasoning over the entire figure (RD). We also explicitly test their abilities to locate relevant panels. Table 3 presents precision, recall and F1 to measure how well models can localize to the correct visual evidence. Their low precision and high recall indicates that models do identify the relevant panel(s), but also deem a lot of irrelevant panels to be important. Clearly, evidence localization is difficult for models.
@@
Claims often encode epistemic relationships which can be systematically perturbed to test the sensitivity of contemporary models. We calculate sensitivity as the percentage of times models change predictions across the supported and refuted version of the same claim. Table 5 shows the sensitivity of models by perturbation type. Models are not sensitive enough to understand nuances in epistemic relationships, being the worst when both the observation and inference is modified. Analyzing differences in models’ confidences for predictions may provide more insight Marcé and Poliak (2022).

### v2

@@ Abstract
Our results show most vision-language models are poor (0.3-0.5 F1), with even the best model only achieving 0.72 F1.
@@ 1 Introduction
However, scientific evidence is often presented as heterogeneous information-rich figures that support the important findings, claims and conclusions of experiments. Therefore, scientific claim verification requires both textual and visual understanding capabilities. To assess a claim, one has to go over the figure and its caption, find the panel(s) with information relevant to the claim, combine this visual knowledge with textual information in the figure caption, and finally judging whether the claim is supported or not. While there is a large number of benchmarks on scientific figures, they focus on image captioning Hsu et al. (2021), question answering Kahou et al. (2017), or other reasoning tasks Yue et al. (2024a). There are no readily usable multimodal benchmarks for scientific claim verification. The closest work, ChartCheck Akhtar et al. (2024), poses a multimodal claim verification task but is restricted to simple data charts crawled from the web, which are substantially different from complex figures found in scientific articles.
To address this gap, we introduce MuSciClaims 111Data is available at https://huggingface.co/datasets/StonyBrookNLP/MuSciClaims/, a multimodal benchmark for claim verification over figures in scientific (physics, chemistry and biology) literature. We set two desiderata for our benchmark: the dataset needs carefully constructed claims that are not supported or have contradictory information in the figures; apart from quantifying model performance, the dataset should also be diagnostic in nature to identify specific model weaknesses. Our dataset creation methodology is designed to meet these desiderata.
We extract claims with inline references to figures from the results section of articles. We manually filter these to only retain claims that are clearly and unambiguously supported by the figures. Then, we create contradictory claims by perturbing these supporting claims. We devise a diverse set of perturbations to test specific capabilities for claim verification including qualitative and quantitative reasoning, and observation-inference connections.
Last, we create a suite of diagnostic tasks associated with each claim to better understand model failures. Specifically, we design tasks that help uncover errors across aspects of basic visual understanding, evidence localization, cross-modal aggregation, and epistemic sensitivity. We ensure the integrity of the dataset through manual analysis. The resulting dataset consists of 1515 (claim, figure) data points from Physics, Chemistry and Biology, equally balanced across 3 class labels (Support, Neutral, Contradict), each accompanied by diagnostic questions.
@@
Their basic visual understanding and cross-modal aggregation capabilities also need improvement.
@@ 2 Related Work
To create more complex QA benchmarks, ChartQA Masry et al. (2022) mixes 30k human and machine-generated questions; however the images are still limited to line, bar and pie charts. To cover more types, ArXivQA Li et al. (2024) extracts images with LLM-generated QA pairs from arXiv papers. SciGraphQA Li and Tajbakhsh (2023) extract graphs from Comp. Sci. ArXiv papers and use LLMs to create multi-turn question-answering dialogues about them. MMC Liu et al. (2024) supports diverse tasks and chart types using free-form questions and open-ended answers.
Previous benchmarks rely heavily on chart annotations or table metadata as textual prompts to generate content, allowing models to easily obtain candidate answers while ignoring the charts’ visual logic. ChartBench Xu et al. (2023) includes both annotated and unannotated charts. While ChartX Xia et al. (2024) covers more chart types, its data and charts are synthesized and limited to ones that can be directly converted into a structural data format, e.g., CSV format. CharXiv Wang et al. (2024) consists of 2k real-world charts with manually curated questions by human experts and answers validated by hand, panning 8 major subjects published on arXiv. The questions are either descriptive to understand basic chart data or reasoning-based to dig deeper into charts. MultiChartQA Zhu et al. (2025) is designed to evaluate VLMs’ reasoning capabilities across multiple charts. However, the charts are not information-rich and no domain knowledge beyond what is stated in the charts is required to answer questions in these benchmarks.
To cover more image types, MMMU Yue et al. (2024a, b) collected multimodal questions from college exams, quizzes, and textbooks, covering six disciplines, ranging from visual scenes like photos and paintings to diagrams and tables, testing the perceptual capabilities of VLMs. CURIE Cui et al. (2025) covers diverse scientific disciplines, but its multimodal tasks are limited to biodiversity georeferencing and protein sequence reconstruction tasks. While this requires domain knowledge, it doesn’t require a high level of expertise.
@@ 3 Creating
MuSciClaims contains a claim, its associated figure and caption and annotations about its relevant panels. Each claim, both original and perturbed, is also labeled with its relationship to the figure (Support, Neutral, Contradict). It also enables performing diagnostic tests. Models must identify the relevant panel as part of EvidenceLocalization and answer a question about the figure for BasicVisualUnderstanding. To perform Cross-ModalAggregation, model performance should drop when given either just the figure or the caption. For EpistemicSensitivity, model predictions must change across a pair of (original, perturbed) claims.
Verifying whether a claim is supported by scientific evidence requires understanding different parts of the claim, locating and extracting information from multimodal sources, reasoning with it, and finally making a judgment.
@@
MuSciClaims, a dataset created from scientific articles across physics, chemistry and biology.
We use open-access, peer-reviewed articles published in Nature Physics, Journal of the American Chemical Society, and Cell, as the source for claims in physics, chemistry and biology respectively.
@@ 3.1 Automatic Extraction
We extract figures and associated claims from the results section of the articles, where key findings and takeaways are described along with supporting evidence, often expressed within heterogeneous figures containing charts, microscopy images, chemical reaction schemes, or other diagrams.
@@ 3.2 Systematic Claim Perturbation
The claims extracted from the articles are grounded in the associated figures—i.e., the figures support222We rely on the scientific integrity of the published articles and assume that evidence support the asserted claims. the claims. To create an effective test bed, we also need claims that are not supported and have contradictory evidence in the figures. Further, we want to ensure that we test for a variety of reasons that could make a claim unsupported or contradictory with respect to the given evidence. To this end, we use a manual claim perturbation process to ensure meaningful perturbations, and ensure their quality through a second annotation process. Annotators produce free-form contradicting claims over a range of perturbations.
We manually analyzed the original claims to identify the main capabilities needed for checking a claim against the corresponding figures. Based on this, we create four categories of perturbations: (i) Qualitative Inference—Directional terms are replaced with their opposites (e.g., “high concentration" to “low concentration"). This tests whether models can check if the asserted qualitative statements are supported via visual relationships between data points in a figure. (ii) Qualitative Relationship Inference—Comparisons are edited (e.g., “X is stronger than Y" into “Y is stronger than X") to create the opposite conclusion from a figure. This checks for assessing qualitative relationships between variables via visual inference of similar relationships, (iii) Quantitative reasoning—Numerical values, primarily associated with experiment details such as statistical significance or experiment size, are modified to test for reasoning about the key quantities of interest. (iv) Epistemic Mismatch—This represents a disconnect between different forms of knowledge. We add perturbations that introduce inconsistencies between an observation that is visually true (i.e., supported by the figure) and its inference (which requires domain knowledge). This tests for the ability to carefully connect visually verified information with the textually asserted effect.
Our perturbations ensure that the modified claim is a contradiction of the supported claim. This means that the figure which supports the original claim will, by extension, not support the modified claim. We verify the quality of the resulting perturbations through a second round of manual annotation. For a subset of data, three annotators were provided a supported claim as well as its perturbation. They are required to judge whether the perturbation contradicts the supported claim. We find that all three annotators agree that all perturbed claims are indeed contradictions of the corresponding supported claims (100% agreement).
@@ 3.4 Dataset Statistics
Through the process described above, we obtain 505 claims that are supported by a figure (Support), and 505 corresponding perturbed claims in contradiction to a given figure (Contradict).
@@
MuSciClaims contains 1515 data points balanced equally across 3 class labels.
Out of this, 918 data points are from biology, 309 from chemistry and 288 from physics.
Each data point is also annotated with figure panels most relevant to a claim, a question about the figure and information about perturbation types to support our diagnostic tests333We will release the data in accordance with the papers’ CC BY 4.0 license..
@@
| 4o-mini | D | 0.41 | 0.88 | 0.56 | 0.64 | 0.48 | 0.55 | 0.75 | 0.10 | 0.17 | 0.60 | 0.48 | 0.43 |
| RD | 0.43 | 0.83 | 0.56 | 0.64 | 0.47 | 0.54 | 0.62 | 0.20 | 0.30 | 0.56 | 0.50 | 0.47 | |
| 4o | D | 0.43 | 0.93 | 0.59 | 0.86 | 0.46 | 0.60 | 0.75 | 0.23 | 0.35 | 0.68 | 0.54 | 0.51 |
| RD | 0.47 | 0.86 | 0.61 | 0.71 | 0.61 | 0.65 | 0.76 | 0.25 | 0.38 | 0.65 | 0.57 | 0.55 | |
| Sonnet | D | 0.52 | 0.87 | 0.65 | 0.83 | 0.64 | 0.72 | 0.79 | 0.43 | 0.56 | 0.71 | 0.65 | 0.64 |
| RD | 0.53 | 0.89 | 0.66 | 0.84 | 0.64 | 0.73 | 0.81 | 0.47 | 0.59 | 0.73 | 0.66 | 0.66 | |
| o3 | RD | 0.67 | 0.74 | 0.71 | 0.69 | 0.79 | 0.74 | 0.82 | 0.61 | 0.70 | 0.73 | 0.72 | 0.72 |
| o4-mini | RD | 0.62 | 0.84 | 0.71 | 0.76 | 0.76 | 0.76 | 0.88 | 0.57 | 0.69 | 0.75 | 0.72 | 0.72 |
| Phi-4 | D | 0.43 | 0.70 | 0.53 | 0.74 | 0.19 | 0.30 | 0.44 | 0.50 | 0.47 | 0.54 | 0.46 | 0.43 |
| RD | 0.36 | 0.81 | 0.51 | 0.81 | 0.10 | 0.17 | 0.58 | 0.26 | 0.36 | 0.58 | 0.41 | 0.34 | |
| Llava-Next | D | 0.37 | 0.92 | 0.53 | 0.68 | 0.33 | 0.44 | 1.00 | 0.00 | 0.01 | 0.68 | 0.42 | 0.33 |
| RD | 0.38 | 0.80 | 0.52 | 0.54 | 0.42 | 0.47 | 0.61 | 0.09 | 0.15 | 0.51 | 0.43 | 0.38 | |
| Llama-3.2 | D | 0.41 | 0.86 | 0.56 | 0.68 | 0.42 | 0.52 | 0.60 | 0.17 | 0.27 | 0.57 | 0.49 | 0.45 |
| RD | 0.37 | 0.93 | 0.53 | 0.72 | 0.15 | 0.25 | 0.61 | 0.17 | 0.27 | 0.56 | 0.42 | 0.35 | |
| Molmo | D | 0.41 | 0.91 | 0.57 | 0.77 | 0.29 | 0.42 | 0.54 | 0.22 | 0.32 | 0.57 | 0.47 | 0.43 |
| RD | 0.39 | 0.75 | 0.51 | 0.56 | 0.28 | 0.37 | 0.43 | 0.23 | 0.30 | 0.46 | 0.42 | 0.39 | |
| InternVL3 | D | 0.62 | 0.79 | 0.70 | 0.83 | 0.68 | 0.75 | 0.70 | 0.64 | 0.67 | 0.72 | 0.70 | 0.70 |
| RD | 0.45 | 0.91 | 0.60 | 0.83 | 0.49 | 0.61 | 0.80 | 0.30 | 0.44 | 0.69 | 0.57 | 0.55 | |
| Qwen2.5 | D | 0.55 | 0.80 | 0.65 | 0.70 | 0.80 | 0.75 | 0.85 | 0.34 | 0.48 | 0.70 | 0.65 | 0.63 |
| RD | 0.42 | 0.91 | 0.57 | 0.79 | 0.41 | 0.54 | 0.85 | 0.25 | 0.39 | 0.68 | 0.53 | 0.50 | |
| DeepSeek | D | 0.55 | 0.43 | 0.48 | 0.50 | 0.67 | 0.58 | 0.44 | 0.40 | 0.42 | 0.50 | 0.67 | 0.58 |
| RD | 0.41 | 0.65 | 0.51 | 0.51 | 0.51 | 0.51 | 0.51 | 0.21 | 0.30 | 0.48 | 0.46 | 0.44 | |
@@ 4.1 Evaluation Tasks
MuSciClaims also supports four diagnostic tasks designed to assess a diverse set of capabilities required to effectively verify claims.
@@
We use Exact Match to judge whether a model answer is correct.
@@
(4) EpistemicSensitivity tests whether models consistently (and correctly) change their prediction across epistemic perturbations of the same claim; they should predict support for the original and contradict for the perturbed claim.
@@
As part of §3.2, annotators also mark whether they perturb the observation, the inference or both.
We perform a sensitivity test for the same across (original, perturbed) claim pairs.
@@ 4.2 Models
We conduct the aforementioned evaluation on a set of different vision-language models (VLMs): gpt-4o-mini-2024-07-18 (4o-mini), gpt-4o-2024-11-20 (4o), claude-3-5-sonnet-20241022 (Sonnet), o3-2025-04-16 (o3), o4-mini-2025-04-16 (o4-mini), Phi-4 Multimodal Instruct (Phi-4), llava-v1.6-mistral-7b-hf (Llava-Next), Llama-3.2-11B Vision Instruct (Llama-3.2), Molmo-7B-D (Molmo), InternVL3-38B (InternVL3), Qwen2.5-VL-32B (Qwen2.5) and deepseek-VL2-small (DeepSeek).
@@
More details are in Appendix A, G and D.
@@ 5 Results
MuSciClaims 444Results for each domain are presented in Table 7, 8 and 9.
@@
We observe that most models perform poorly on the task (D rows in Table 1), with overall F1 scores only ranging from 0.3-0.5. Only two models (out of ten) stand out: Sonnet (0.66 F1) and InternVL3 (0.70 F1). Models attain high recall and low precision on Support. In contrast for the Neutral and Contradict claims, models have low recall and high precision. This implies two findings: First, models have a strong bias towards recognizing most of the claims as supported. Second, the models can reliably identify some of the Neutral and Contradict claims. Our manual analysis shows that models only identify the most obviously wrong claims as the Neutral and Contradict claims. This can be done reliably but they struggle on ones that are more difficult, which require careful reasoning. These suggest challenges for claim verification methods.
@@
The RD rows in Table 1 show results where models, given the figure and caption, first perform step-by-step reasoning on the claim and then generate their decision on the category of the claim. Results show that reasoning leads to improvements (0.02-0.04) for closed-source models and Llava-Next, but the gains are rather small. o3 and o4-mini, models trained to analyze and do reasoning over images, achieve the highest performance (0.72 F1).
@@
Table 6 presents model performance for different domains in MuSciClaims. On average, models are worst at verifying Physics claims and best at judging Chemistry claims. However, the highest performance is achieved on Biology claims.
@@ 6 Diagnostics Results
We use our diagnostic tests to better understand the failure modes of 4o-mini, 4o, Sonnet and InternVL3.
Going forward, we run these diagnostic tests (§3.3) on the Biology subset of
MuSciClaims and discuss them.
@@
Finding the most relevant panels of figures is important to assess claims from information-rich figures. Table 2 shows how well models perform when prompted to first identify the associated panels, reason over them and make a decision (IRD), thereby testing EvidenceLocalization. Model performance deteriorates when localizing before reasoning (IRD) as compared to reasoning over the entire figure (RD). We also explicitly test their ability to locate relevant panels. Table 3 presents precision, recall and F1 to measure how well models can localize to the correct visual evidence. Their low precision and high recall indicates they do identify the relevant panel(s), but also deem a lot of irrelevant panels to be important. Clearly, evidence localization is difficult for models.
@@
Claims often encode epistemic relationships which can be systematically perturbed to test the sensitivity of contemporary models Verma et al. (2023). We calculate sensitivity as the percentage of times models change predictions across the supported and refuted version of the same claim. Table 5 shows the sensitivity of models by perturbation type. Models are not sensitive enough to understand nuances in epistemic relationships, being the worst when both the observation and inference is modified. Analyzing differences in models’ confidences for predictions may provide more insight Marcé and Poliak (2022).